OpenStudy (anonymous):

sqrt(48+sqrt(x))-4=4thsqrt(x) find x?

OpenStudy (anonymous):

\[\sqrt{48+\sqrt{x}}-4=\sqrt[4]{x}\]

OpenStudy (anonymous):

\[\sqrt{48+\sqrt{x}}-4=\sqrt[4]{x}\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh yes, i see it

OpenStudy (anonymous):

guess you this is going to be annoying. you are going to have to square both sides then do it again

OpenStudy (anonymous):

i don't get it ?

OpenStudy (anonymous):

yeah i can see this is going to be a pain square both sides \[(\sqrt{48+\sqrt{x}}-4)^2=(\sqrt[4]{x})^2\]

OpenStudy (anonymous):

ooh i see, and then i do another on

OpenStudy (anonymous):

you get \[48+\sqrt{x}-8\sqrt{48+\sqrt{x}}+16\sqrt{x}\]

OpenStudy (anonymous):

actually you get \[48+\sqrt{x}-8\sqrt{48+\sqrt{x}}+16=\sqrt{x}\]

OpenStudy (anonymous):

we can cancel the \(\sqrt{x}\) from both sides also \(48+16=64\) so we get \[64-8\sqrt{48+\sqrt{x}}=0\]

OpenStudy (anonymous):

subtract 64 from both sides divide both sides by -8 and get \[\sqrt{48+\sqrt{x}}=8\]

OpenStudy (anonymous):

square again, and you should be in good shape don't forget to check your answer because squaring could introduce and extraneous solution